# OpenSlot: Mixed Open-set Recognition with Object-centric Learning

### v1

@@ VI Conclusion
This paper introduced the mixed open-set recognition (OSR) problem, which considers the joint occurrence of unknown and known class objects in negative images. To tackle this challenge, we presented the OpenSlot framework, built upon the object-centric learning approach. Through the proposed anti-noise-slot (ANS) technique, we effectively excluded noisy slot features from the image classification process, addressing the semantic misalignment issue inherent in slot-based class predictions. Extensive experiments and ablation studies demonstrated OpenSlot’s superiority in handling the mixed OSR problem, significantly outperforming existing OSR studies and achieving SOTA results in conventional OSR benchmarks. Moreover, OpenSlot can be extended to open-set object detection to localize unknown class objects and explicitly explain label shifts. Competitive results and advantages in computational efficiency highlight the benefits of our model.
@@
VOC-6/142400270964170938017461361
COCO-20/60800012892619671962514781932013132
VOC-COCO5717109912713533635120632114852335
Task1 (20/60)98124877311913834886903196643412
Task2 (40/40)21235104051686817583274111963253063
Task3 (60/20)46671231247845413129125943034329
TABLE X: Summary of the single/multi-label OSR tasks. The values are the number of images in the individual datasets.
In this appendix, we introduce the detailed benchmark information and implementations.
-A Mixed OSR
For single-label tasks, we use all KKC and UUC candidates to construct VOC-6/14 and COCO-20/60. We use VOC-Cls and COCO to conduct the multi-label experiments and build up the test sets with the same criteria as the single-label setting. In COCO-20/60, 40/40, 60/20 tasks, KKCs are randomly sampled, and the results vary with the sampling effects. We report the details of all OSR tasks in Tab. VI:
-B Experiment details
Open-set object detection. We use the noise classifier $\varphi_{nz}$’s predictions and a threshold (0.75) to determine the foreground (FG) slots (below threshold). We then mask (over 0.5) the pixels in the FG attention mask and consider the resulting binary mask as the slot’s semantic region. Assuming a meaningful semantic object must be over 40 pixels, we obtain a valid object mask to compute bounding boxes. Finally, we score [27] all FG slots and consider those with scores over 5 as known objects and the rest as unknown.

### v2

@@ VI Conclusion
This paper introduces the mixed open-set recognition (OSR) problem, which considers the joint occurrence of known and unknown class objects in negative images. To address this challenge, we present the OpenSlot framework, built upon the object-centric learning approach. Through the proposed anti-noise-slot (ANS) technique, we effectively distinguish noisy slots from true class features and exclude them from the classification process, thus addressing the semantic misalignment issue inherent in slot-based class predictions. Extensive experiments and ablation studies demonstrated that OpenSlot significantly outperforms existing studies in handling the mixed OSR problem and achieves SOTA results in conventional OSR benchmarks. Additionally, OpenSlot can be extended to open-set object detection, localizing unknown class objects and explicitly explaining label shifts. The competitive results and computational efficiency show the advantages of our model. In this appendix, we introduce the detailed benchmark information and implementations.
-A Mixed OSR
For single-label tasks, we use all KKC and UUC candidates to construct VOC-6/14 and COCO-20/60. We use VOC-Cls and COCO to conduct the multi-label experiments and build up the test sets with the same criteria as the single-label setting. In COCO-20/60, 40/40, 60/20 tasks, KKCs are randomly sampled, and the results vary with the sampling effects.
-B Experiment details
Open-set object detection. We use the noise classifier $\varphi_{nz}$’s predictions and a threshold (0.75) to determine the foreground (FG) slots (below threshold). We then mask (over 0.5) the pixels in the FG attention mask and consider the resulting binary mask as the slot’s semantic region. Assuming a meaningful semantic object must be over 40 pixels, we obtain a valid object mask to compute bounding boxes. Finally, we score [29] all FG slots and consider those with scores over five as known objects and the rest as unknown.
@@
[1pt]-
30.843.956.435.526.014.066.553.832.67.136.716.236.738.752.58.212.736.242.955.950.9
[…]
TABLE X: Category-level $AP_{50}$ results on Pascal Voc 2007.